### v1

@@ 6.2 $\epsilon$-Regularizer
Suppose that $c\geq 1$ and $d\leq 1$, and let $\eta:=2/(2+d)$. Let $E=\Theta(|\mathcal{A}|^{-\eta}(|\mathcal{B}|\log T)^{1-\eta}(c\cdot T)^{\eta})$. Let $\texttt{ALG}_{2}$ be any algorithm with high-probability instantaneous regret
$g(t,T,\mathcal{B})=O\left((|\mathcal{A}|\cdot|\mathcal{B}|\cdot\log T)^{\eta/2}\cdot(c\cdot T)^{-\eta/2}\right)$ for $t>E$ and $g(t,T,\mathcal{B})=1$ for $t\leq E$, and let $\texttt{ALG}_{1}=\texttt{ExploreThenUCB}(E)$. Then, then the leader and follower regret with respect to the generalized $(c,d,\gamma)$-tolerant benchmarks are both bounded as:
@@
The proofs of Theorem 6.2 and Theorem 6.3 follows from the same arguments as the proof of Theorem 4.4 and Theorem 4.5, respectively, but with the values of $E_{1},E_{2}$ modified (full proofs are deferred to Appendix F). Note that as $d$ decreases, the regret bound worsens: this aligns with the intuition that smaller values of $d$ capture greater intolerance. Similarly, the regret increases with $c$.
@@
7 Algorithms satisfying high-probability instantaneous regret and high-probability anytime regret bounds
Our algorithms for the leader placed assumptions on the fine-grained performance of the follower’s algorithm. More specifically, the regret bound for ExploreThenUCB required an high-probability instantaneous regret bound for the follower (Theorem 4.5), and the regret bound for LipschitzUCB required an high-probability anytime regret bound for the follower (Theorem 5.1). In this section, we discuss these conditions in more detail and analyze which standard algorithms satisfy each of these conditions.
First, we observe that high probability instantaneous regret bound immediately translate to high-probability anytime regret bounds.
@@
$h(t,T):=\sum_{t^{\prime}=1}^{t}g(t^{\prime},T)=E+\sum_{t^{\prime}=E}^{t}O\left((|\mathcal{A}||\mathcal{B}|\log T)^{1/3}T^{-1/3}\right)=O\left((|\mathcal{A}||\mathcal{B}|\log T)^{1/3}T^{2/3}\right).$
However, this naive high-probability anytime regret bound is not strong enough for Theorem 5.1: we can nonetheless achieve the desired regret bound with additional assumptions on $\texttt{ALG}_{2}$ as we describe below.
Initialize $s^{\prime}=0$, $t^{\prime}=1$, $\mathcal{B}^{\prime}=\mathcal{B}$   // Index of the last completed phase, time step marking beginning of phase $s^{\prime}+1$, active arms in phase $s^{\prime}$.
Initialize $newphase=\text{False}$.   // Boolean for first time step in phase.
@@
                   Set $S(a,b):=\left\{r\mid\exists(t^{\prime\prime},a_{t^{\prime\prime}},b_{t^{\prime\prime}},r)\in H\text{ s.t. }a_{t^{\prime\prime}}=a,b=b_{t^{\prime\prime}},t^{\prime\prime}\geq t^{\prime}\right\}$   // observed rewards
                   $\hat{v}_{2}(a,b)\leftarrow(\sum_{r\in S(a,b)}r)/|S(a,b)|$   // compute empirical mean
@@
$i$ = $((t-t^{\prime})\mod(\left|\mathcal{B}^{\prime}\right|))+1$// Calculate next arm to be pulled
@@
Next, we show that ActiveArmElimination (Even-Dar et al., 2002) (Algorithm 7, see
Lattimore and Szepesvári (2020) for a textbook treatment) satisfies both the high-probability instantaneous regret bound required for Theorem 4.5 and the high-probability anytime regret bound required for Theorem 5.1.
@@
Note that ExploreThenCommit does not satisfy the high-probability anytime regret bound required for Theorem 5.1 due to the uniform exploration phase at the beginning of the algorithm means.
@@ 9 Acknowledgements
Appendix A Relaxed benchmarks in example instances
@@ A.3 Additional worked out example for the benchmark
Appendix B Additional Notation and Auxiliary Lemmas
@@
In the weakly decentralized setting, let the leader’s history up to time step $t$ be the set of arms that were pulled, as well as the reward for the leader at each time step:
@@
In the strongly decentralized setting, the leader cannot even observe the action chosen by the follower, but the follower’s information remains unchanged. That is $H_{1,t}:=\left\{(t^{\prime},a_{t^{\prime}},r_{1,t^{\prime}}(a_{t^{\prime}},b_{t^{\prime}}))\mid 1\leq t^{\prime}<t\right\}$.
@@
Appendix C Proofs of regret lower bounds
Our regret bounds analyze a centralized setting (Section C.1) and build on standard tools [Lattimore and Szepesvári, 2020] for regret lower bounds (Section C.2). We prove Proposition 5.4 in Section C.3, Theorem 3.1 in Section C.4, and Theorem 4.6 in Section C.5.
C.1 Centralized environment
@@
Lemma C.1.
Fix an instance $\mathcal{I}=(\mathcal{A},\mathcal{B},v_{1},v_{2})$ and time horizon $T$. Let $\texttt{ALG}_{1}$, $\texttt{ALG}_{2}$ be any pair of decentralized algorithms operating in either the strongly decentralized setting or weakly decentralized setting. Then, there exists a centralized algorithm ALG such that the leader rewards $(r_{1,1}(a_{1},b_{1}),\ldots,r_{1,T}(a_{T},b_{T}))$ are identically distributed for ALG and $(\texttt{ALG}_{1},\texttt{ALG}_{2})$ and the follower rewards $(r_{2,1}(a_{1},b_{1}),\ldots,r_{2,T}(a_{T},b_{T}))$ are also identically distributed for ALG and $(\texttt{ALG}_{1},\texttt{ALG}_{2})$.
Lemma C.1 follows immediately from designing ALG to “simulate” histories for the leader and the follower (by projecting away the information unavailable to each player) and then to choose arms by applying $\texttt{ALG}_{1}$ and $\texttt{ALG}_{2}$ on these histories.
C.2 Useful lemmas
@@
Theorem C.2 (paraphrased from Theorem 14.2 in [Lattimore and Szepesvári, 2020]).
@@
Theorem C.3 (adapted from Lemma 15.1 in [Lattimore and Szepesvári, 2020]).
@@
Recall that we assume Gaussian noise, which further simplifies Theorem C.3. By applying standard KL divergence bounds for univariate Gaussians, we obtain the following corollary of Theorem C.3.
Corollary C.4.
@@
Table 7: Hard instance for Proposition 5.4, where *** is equal to $(0,0)$ for instance $\mathcal{I}_{a_{1},b_{1}}$, and $(2\delta,2\delta)$ otherwise.
@@
We define a family of instances in the centralized game and evaluate the self-tolerant benchmarks on this family of instances. Arbitrarily pick some $a_{1}\in\mathcal{A}$ to be the “base” action. Let $\mathcal{F}_{\delta,\mathcal{A},\mathcal{B}}$ be the family of $(|\mathcal{A}|-1)\cdot|\mathcal{B}|+1$ instances of the form $(\mathcal{A},\mathcal{B},v_{1},v_{2})$ for varying settings of $v_{1}$ and $v_{2}$, where we index the instances by $(a^{\prime},b^{\prime})\in((\mathcal{A}\setminus\left\{a_{1}\right\})\times\mathcal{B})\cup\left\{(a_{1},b_{1})\right\}$. The utility functions for the instance $\mathcal{I}_{(a^{\prime},b^{\prime})}$ are equal to the terms below (illustrated in Table 7):
@@
Fix any ALG for the centralized environment. For each $(a,b)\in((\mathcal{A}\setminus\left\{a_{1}\right\})\times\mathcal{B})\cup\left\{(a_{1},b_{1})\right\}$, let $P_{a,b}$ denote the probability measure over canonical centralized bandit model when ALG is applied to the instance $\mathcal{I}_{a,b}$ (see Section C.2). Let $n_{T}(a,b)=\sum_{t=1}^{T}\mathbbm{1}[a_{t}=a,b_{t}=b]$ be the random variable denoting the number of times that $(a,b)$ is pulled. We define:
@@
where $G^{c}$ is the complement of $G$. We apply Theorem C.2 to see that:
@@
where (1) applies Theorem C.2 and (2) applies Corollary C.4, and (3) applies the fact that $n_{T}(a_{m},b_{m})\leq\frac{T}{(|\mathcal{A}|-1)\cdot|\mathcal{B}|}$. If we set $\delta=\Theta(\sqrt{\frac{|\mathcal{A}-1)|\mathcal{B}|}{T}})$, then we obtain a bound of $\Theta(\sqrt{T\cdot(|\mathcal{A}|-1)\cdot|\mathcal{B}|})$. Since expected regret is nonnegative for these instances (see discussion above), this implies that either $R(T;\mathcal{I}_{a_{1},b_{1}})=\Omega(\sqrt{T\cdot(|\mathcal{A}|-1)\cdot|\mathcal{B}|})$ or $R(T;\mathcal{I}_{a_{m},b_{m}})=\Omega(\sqrt{T\cdot(|\mathcal{A}|-1)\cdot|\mathcal{B}|})$ as desired.
C.4 Proof of Theorem 3.1
@@
It suffices to prove this lower bound in a centralized environment where a single learner can choose action pairs $(a,b)$ and observes rewards for both players (Lemma C.1). We construct a pair of instances $\mathcal{I}$ and $\tilde{\mathcal{I}}$ such that at least one of the players incurs linear regret on at least one of the instances. In particular, we take $\mathcal{I}$ and $\tilde{\mathcal{I}}$ to be the instances depicted in Table 1 with $\delta=O(1/\sqrt{T})$ (reproduced here for convenience).
@@
Fix any ALG for the centralized environment. Let $P$ (resp. $\tilde{P}$) denote the probability measure over canonical centralized bandit model when ALG is applied to the instance $\mathcal{I}$ (resp. $\tilde{\mathcal{I}}$) (see Section C.2). We will show that the regret is high on at least one of the instances $\mathcal{I}$ and $\tilde{\mathcal{I}}$.
@@
where $G^{c}$ is the complement of $G$. We apply Theorem C.2 to see that:
@@
where (1) applies Theorem C.2 and (2) applies Corollary C.4, and (3) uses the fact that $n_{T}(a_{1},b_{2})\leq T$. If we take $\delta=O(T^{-1/2})$, then we obtain a bound of $\Omega(T)$. Since these expected regrets are always nonnegative (see discussion above), this implies that either $R_{1}(T;\mathcal{I})=\Omega(T)$ or $R_{2}(T;\tilde{\mathcal{I}})=\Omega(T)$ as desired. ∎
C.5 Proof of Theorem 4.6
@@
Table 8: Hard instance for Theorem 4.6, where *** is equal to $(0,0)$ for instance $\mathcal{I}_{a_{1},b_{1}}$, and $(0,2\delta)$ otherwise. Note that this example is structurally similar to the illustrative example in Table 3, but with $\left|\mathcal{A}\right|,\left|\mathcal{B}\right|\geq 2$.
@@
We define a family of instances in the centralized game and evaluate the self-tolerant benchmarks on this family of instances. Arbitrarily pick some $(a_{1},b_{1})\in\mathcal{A}\times\mathcal{B}$ to be the “base” action. Let $\mathcal{F}_{\delta,\mathcal{A},\mathcal{B}}$ be the family of $|\mathcal{B}|$ instances of the form $(\mathcal{A},\mathcal{B},v_{1},v_{2})$ for varying settings of $v$, where we index the instances by $\mathcal{B}$. The utility functions for the instance $\mathcal{I}_{b^{\prime}}$ are equal to terms below (illustrated in Table 8):
@@
Fix an ALG be an algorithm for the centralized environment. For each $b\in\mathcal{B}$, let $P_{b}$ denote the probability measure over canonical centralized bandit model when ALG is applied to the instance $\mathcal{I}_{b}$ (see Section C.2). Let $n_{T}(a,b)=\sum_{t=1}^{T}\mathbbm{1}[a_{t}=a,b_{t}=b]$ be the random variable denoting the number of times that $(a,b)$ is pulled. We define:
@@
where $G^{c}$ is the complement of $G$. We apply Theorem C.2 to see that:
@@
where (1) applies Theorem C.2 and (2) applies Corollary C.4 and where (3) uses the fact that $\sum_{b\neq b_{1}}\sum_{a\neq a_{1}}\mathbb{E}_{P}[n_{T}(a,b)]\geq(|\mathcal{B}|-1)\sum_{a\neq a_{1}}\mathbb{E}_{P}[n_{T}(a,b_{m})]$.
@@
Appendix D Proofs for Section 4
D.1 Proof of Proposition 4.3

### v2

@@ 6.2 $\epsilon$-Regularizer
Suppose that $c\geq 1$ and $d\leq 1$, and let $\eta:=2/(2+d)$. Let $E=\Theta(|\mathcal{A}|^{-\eta}(|\mathcal{B}|\log T)^{1-\eta}(c\cdot T)^{\eta})$. Consider a StrongDSG where $\texttt{ALG}_{2}$ is any algorithm with high-probability instantaneous regret
$g(t,T,\mathcal{B})=O\left((|\mathcal{A}|\cdot|\mathcal{B}|\cdot\log T)^{\eta/2}\cdot(c\cdot T)^{-\eta/2}\right)$ for $t>E$ and $g(t,T,\mathcal{B})=1$ for $t\leq E$, and where $\texttt{ALG}_{1}=\texttt{ExploreThenUCB}(E)$. Then, then the leader and follower regret with respect to the generalized $(c,d,\gamma)$-tolerant benchmarks are both bounded as:
@@
The proofs of Theorem 6.2 and Theorem 6.3 follows from the same arguments as the proof of Theorem 4.4 and Theorem 4.5, respectively, but with the values of $E_{1},E_{2}$ modified (full proofs are deferred to Appendix E). Note that as $d$ decreases, the regret bound worsens: this aligns with the intuition that smaller values of $d$ capture greater intolerance. Similarly, the regret increases with $c$.
@@
7 Discussion of Assumptions on the Follower’s Algorithm
Our algorithms for the leader placed assumptions on the fine-grained performance of the follower’s algorithm. More specifically, the regret bound for ExploreThenUCB required an high-probability instantaneous regret bound $g$ for the follower (Theorem 4.5), and the regret bound for LipschitzUCB required an high-probability anytime regret bound $h$ for the follower (Theorem 5.1).
In this section, we examine these two conditions in more detail. First, we relate these two conditions and show that many standard algorithms satisfy the conditions on $g$ and $h$ in Theorem 4.5 and Theorem 5.1 (Section 7.1). Then, we extend our analysis of ExploreThenUCB and LipschitzUCB to more general conditions on $g$ and $h$, respectively (Section 7.2).
7.1 Algorithms Satisfying These Fine-Grained Regret Guarantees
As a warmup, we first observe that high probability instantaneous regret bounds immediately translate to high-probability anytime regret bounds.
@@ 7.1 Algorithms Satisfying These Fine-Grained Regret Guarantees
$h(t,T):=\sum_{t^{\prime}=1}^{t}g(t^{\prime},T)=E+\sum_{t^{\prime}=E+1}^{t}O\left((|\mathcal{A}||\mathcal{B}|\log T)^{1/3}T^{-1/3}\right)=O\left((|\mathcal{A}||\mathcal{B}|\log T)^{1/3}T^{2/3}\right).$
However, this naive high-probability anytime regret bound is not strong enough for Theorem 5.1. We can nonetheless achieve the desired regret bound with additional assumptions on $\texttt{ALG}_{2}$ as we describe below.
Initialize $s^{\prime}=0$, $t^{\prime}=1$, $\mathcal{B}^{\prime}=\mathcal{B}$
  // Index of the last completed phase, time step marking beginning of phase $s^{\prime}+1$, active arms in phase $s^{\prime}$.
Initialize $newphase=\text{False}$.
  // Boolean for first time step in phase.
@@
                   Set $S(a,b):=\left\{r\mid\exists(t^{\prime\prime},a_{t^{\prime\prime}},b_{t^{\prime\prime}},r)\in H\text{ s.t. }a_{t^{\prime\prime}}=a,b=b_{t^{\prime\prime}},t^{\prime\prime}\geq t^{\prime}\right\}$
                    // observed rewards
                   $\hat{v}_{2}(a,b)\leftarrow(\sum_{r\in S(a,b)}r)/|S(a,b)|$
                    // compute empirical mean
@@
$i$ = $((t-t^{\prime})\mod(\left|\mathcal{B}^{\prime}\right|))+1$.
  // Calculate next arm to be pulled
@@
By leveraging the structural properties of specific algorithms, we show that many standard algorithms achieve high-probability instantaneous regret $g$ and/or high-probability anytime regret $h$, where $g$ and $h$ are specified according to the functional forms in Theorem 4.5 and Theorem 5.1. Proofs of these results are deferred to Section F.1.
First, we show that ActiveArmElimination (Even-Dar et al., 2002) (Algorithm 7,
see Lattimore and Szepesvári (2020) for a textbook treatment) satisfies both the high-probability instantaneous regret bound required for Theorem 4.5 and the high-probability anytime regret bound required for Theorem 5.1.
@@
Note that ExploreThenCommit does not satisfy the high-probability anytime regret bound required for Theorem 5.1 due to the uniform exploration phase at the beginning of the algorithm.
@@
7.2 Generalized Analysis of ExploreThenUCB and LipschitzUCB
While the specific instantations of $g$ and $h$ in Theorem 4.5 and Theorem 5.1 are tailored to standard algorithms (Section 7.1), we generalize our analysis of ExploreThenUCB and LipschitzUCB to a richer class of functions $g$ and $h$, respectively.
We generalize Theorem 4.5 to functions $g(t,T)=O(E^{-c_{1}}|\mathcal{B}|^{c_{2}}(\log T)^{c_{3}})$ for $t>E$, where $c_{1}\in(0,1)$ and $c_{2},c_{3}>0$ are arbitrary parameters and where $E=\Theta(|\mathcal{A}|^{-1/(1+c_{1})}|\mathcal{B}|^{c_{2}/(1+c_{1})}\log(T)^{c_{2}/(1+c_{1})}\cdot T^{1/(1+c_{1})})$.
Theorem 7.4.
Let $c_{1}\in(0,1)$ and $c_{2},c_{3}>0$. Let $E=\Theta(|\mathcal{A}|^{-1/(1+c_{1})}|\mathcal{B}|^{c_{2}/(1+c_{1})}(\log T)^{c_{3}/(1+c_{1})}\cdot T^{1/(1+c_{1})})$. Consider a StrongDSG where $\texttt{ALG}_{2}$ is any algorithm with high-probability instantaneous regret $g(t,T,\mathcal{B})=O\left(E^{-c_{1}}|\mathcal{B}|^{c_{2}}(\log T)^{c_{3}}\right)$ for $t>E$ and $g(t,T,\mathcal{B})=1$ for $t\leq E$, and where $\texttt{ALG}_{1}=\texttt{ExploreThenUCB}(E)$. Then, it holds that the regret with respect to the $\gamma$-tolerant benchmarks $\beta_{1}^{\text{tol}}$ and $\beta_{2}^{\text{tol}}$ is bounded as:
$\max(R_{1}(T),R_{2}(T))=O\left(T^{1/(1+c_{1})}\cdot|\mathcal{A}|^{c_{1}/(1+c_{1})}\cdot|\mathcal{B}|^{c_{2}/(1+c_{1})}\cdot(\log T)^{c_{3}/(1+c_{1})}\right)+\Theta\left(\sqrt{T|\mathcal{A}|\log T}\right).$
Note that the special case of $c_{1}=c_{2}=c_{3}=1/2$ recovers the functional form of $g$ in Theorem 4.5. The proof, which follows similarly to the proof of Theorem 4.5, is deferred to Appendix F.2.
We similarly generalize Theorem 5.1 to functions $h(t,T)=C^{\prime}\cdot t^{c_{1}}\cdot|\mathcal{B}|^{c_{2}}\cdot(\log(T))^{c_{3}}$ for $t>E$, where $c_{1},c_{2},c_{3}\in(0,1)$ are arbitrary parameters. This result requires the leader to instead run LipschitzUCBGen (Algorithm 8), a generalized version of LipschitzUCB which adjusts the confidence set size based on the parameters $c_{1}$, $c_{2}$, and $c_{3}$.
Theorem 7.5.
Let $c_{1}\in(0,1)$, $c_{2},c_{3}>0$, and $C^{\prime}>0$ be arbitrary constants. Consider a StrongDSG where $\mathcal{I}=(\mathcal{A},\mathcal{B},v_{1},v_{2})$ has Lipschitz constant $L^{*}$. Let $\texttt{ALG}_{2}$ be any algorithm satisfying high-probability anytime regret $h(t,T,\mathcal{B})=C^{\prime}\cdot t^{c_{1}}\cdot|\mathcal{B}|^{c_{2}}\cdot(\log(T))^{c_{3}}$. Let $\texttt{ALG}_{1}=\texttt{LipschitzUCBGen}(L,C^{\prime}B^{c_{2}},c_{1},c_{3})$ for any $L\geq L^{*}$. Then both players achieve the following regret bounds with respect to the original Stackelberg benchmarks $\beta_{1}^{\text{orig}}$ and $\beta_{2}^{\text{orig}}$: that is, $R_{1}(T;\mathcal{I})=O\left(\sqrt{T|\mathcal{A}||\mathcal{B}|\log T}+L|\mathcal{A}|^{1-c_{1}}|\mathcal{B}|^{c_{2}}(\log T)^{c_{3}}T^{c_{1}}\right)$ and $R_{2}(T;\mathcal{I})=O\left(L\sqrt{T|\mathcal{A}||\mathcal{B}|\log T}+L^{2}|\mathcal{A}|^{1-c_{1}}|\mathcal{B}|^{c_{2}}T^{c_{1}}(\log T)^{c_{3}}\right)$.
Again, note that the special case of $c_{1}=c_{2}=c_{3}=1/2$ recovers the functional form of $g$ in Theorem 5.1. The proof, which follows similarly to the proof of Theorem 5.1, is deferred to Appendix F.2.
Initialize $\hat{v}_{1}(a)=1$ for $a\in\mathcal{A}$.
  // Initialize empirical means for $\max_{b\in\mathcal{B}}v_{1}(a,b)$.
Initialize $v^{\text{UCB}}_{1}(a)=1$ for $a\in\mathcal{A}$.
  // Initialize UCB for $\max_{b\in\mathcal{B}}v_{1}(a,b)$
1 for $a\in\mathcal{A}$ do
       Set $S(a):=\left\{r\mid\exists(t^{\prime},a_{t^{\prime}},r)\in H\text{ s.t. }a=a_{t^{\prime}}\right\}$
        // Observed rewards
2       if $S(a)\neq\emptyset$ then
             $\hat{v}_{1}(a)\leftarrow(\sum_{r\in S(a)}r)/|S(a)|$
              // Empirical mean
             $\alpha(a)\leftarrow\frac{10\sqrt{\mathcal{B}\log T}}{\sqrt{|S(a)|}}+C\cdot L\cdot(\log T)^{c_{3}}T^{c_{1}-1}$
              // confidence bound width
3             $v^{\text{UCB}}_{1}(a)\leftarrow\min(1,\hat{v}_{1}(a)+\alpha(a))$
Let $a^{*}=\text{argmax}_{a\in\mathcal{A}}\left(v^{\text{UCB}}_{1}(a)\right)$.
  // arm with max upper confidence bound
return point mass at $a^{*}$.
Algorithm 8 $\texttt{LipschitzUCBGen}(L,C,c_{1},c_{3})$ applied to $H$
@@ Acknowledgements
Appendix
[…]
Appendix A Worked-out examples, auxiliary notation, and auxiliary lemmas
@@ A.3 Additional worked out example for the benchmark
A.4 Additional Notation and Auxiliary Lemmas
@@
In a weakly decentralized Stackelberg game (WeakDSG), let the leader’s history up to time step $t$ be the set of arms that were pulled, as well as the reward for the leader at each time step:
@@
In a strongly decentralized Stackelberg game (StrongDSG), the leader cannot even observe the action chosen by the follower, but the follower’s information remains unchanged. That is $H_{1,t}:=\left\{(t^{\prime},a_{t^{\prime}},r_{1,t^{\prime}}(a_{t^{\prime}},b_{t^{\prime}}))\mid 1\leq t^{\prime}<t\right\}$.
@@
Appendix B Proofs of regret lower bounds
Our regret bounds analyze a centralized setting (Section B.1) and build on standard tools [Lattimore and Szepesvári, 2020] for regret lower bounds (Section B.2). We prove Proposition 5.4 in Section B.3, Theorem 3.1 in Section B.4, and Theorem 4.6 in Section B.5.
B.1 Centralized environment
@@
Lemma B.1.
Consider a StrongDSG or WeakDSG. Fix an instance $\mathcal{I}=(\mathcal{A},\mathcal{B},v_{1},v_{2})$ and time horizon $T$. For any pair of algorithms $\texttt{ALG}_{1}$ and $\texttt{ALG}_{2}$, there exists a centralized algorithm ALG such that the leader rewards $(r_{1,1}(a_{1},b_{1}),\ldots,r_{1,T}(a_{T},b_{T}))$ are identically distributed for ALG and $(\texttt{ALG}_{1},\texttt{ALG}_{2})$ and the follower rewards $(r_{2,1}(a_{1},b_{1}),\ldots,r_{2,T}(a_{T},b_{T}))$ are also identically distributed for ALG and $(\texttt{ALG}_{1},\texttt{ALG}_{2})$.
Lemma B.1 follows immediately from designing ALG to “simulate” histories for the leader and the follower (by projecting away the information unavailable to each player) and then to choose arms by applying $\texttt{ALG}_{1}$ and $\texttt{ALG}_{2}$ on these histories.
B.2 Useful lemmas
@@
Theorem B.2 (paraphrased from Theorem 14.2 in [Lattimore and Szepesvári, 2020]).
@@
Theorem B.3 (adapted from Lemma 15.1 in [Lattimore and Szepesvári, 2020]).
@@
Recall that we assume Gaussian noise, which further simplifies Theorem B.3. By applying standard KL divergence bounds for univariate Gaussians, we obtain the following corollary of Theorem B.3.
Corollary B.4.
@@
Table 7: Hard instance for Proposition 5.4, where * is equal to $(0,0)$ for instance $\mathcal{I}_{a_{1},b_{1}}$, and $(2\delta,2\delta)$ otherwise.
@@
It suffices to prove this lower bound in a centralized environment where a single learner can choose action pairs $(a,b)$ and observes rewards for both players (Lemma B.1). We define a family of instances in the centralized game and evaluate the self-tolerant benchmarks on this family of instances. Arbitrarily pick some $a_{1}\in\mathcal{A}$ to be the “base” action. Let $\mathcal{F}_{\delta,\mathcal{A},\mathcal{B}}$ be the family of $(|\mathcal{A}|-1)\cdot|\mathcal{B}|+1$ instances of the form $(\mathcal{A},\mathcal{B},v_{1},v_{2})$ for varying settings of $v_{1}$ and $v_{2}$, where we index the instances by $(a^{\prime},b^{\prime})\in((\mathcal{A}\setminus\left\{a_{1}\right\})\times\mathcal{B})\cup\left\{(a_{1},b_{1})\right\}$. The utility functions for the instance $\mathcal{I}_{(a^{\prime},b^{\prime})}$ are equal to the terms below (illustrated in Table 7):
@@
Fix any ALG for the centralized environment. For each $(a,b)\in((\mathcal{A}\setminus\left\{a_{1}\right\})\times\mathcal{B})\cup\left\{(a_{1},b_{1})\right\}$, let $P_{a,b}$ denote the probability measure over canonical centralized bandit model when ALG is applied to the instance $\mathcal{I}_{a,b}$ (see Section B.2). Let $n_{T}(a,b)=\sum_{t=1}^{T}\mathbbm{1}[a_{t}=a,b_{t}=b]$ be the random variable denoting the number of times that $(a,b)$ is pulled. We define:
@@
where $G^{c}$ is the complement of $G$. We apply Theorem B.2 to see that:
@@
where (1) applies Theorem B.2 and (2) applies Corollary B.4, and (3) applies the fact that $n_{T}(a_{m},b_{m})\leq\frac{T}{(|\mathcal{A}|-1)\cdot|\mathcal{B}|}$. If we set $\delta=\Theta(\sqrt{\frac{|\mathcal{A}-1)|\mathcal{B}|}{T}})$, then we obtain a bound of $\Theta(\sqrt{T\cdot(|\mathcal{A}|-1)\cdot|\mathcal{B}|})$. Since expected regret is nonnegative for these instances (see discussion above), this implies that either $R(T;\mathcal{I}_{a_{1},b_{1}})=\Omega(\sqrt{T\cdot(|\mathcal{A}|-1)\cdot|\mathcal{B}|})$ or $R(T;\mathcal{I}_{a_{m},b_{m}})=\Omega(\sqrt{T\cdot(|\mathcal{A}|-1)\cdot|\mathcal{B}|})$ as desired.
B.4 Proof of Theorem 3.1
@@
It suffices to prove this lower bound in a centralized environment where a single learner can choose action pairs $(a,b)$ and observes rewards for both players (Lemma B.1). We construct a pair of instances $\mathcal{I}$ and $\tilde{\mathcal{I}}$ such that at least one of the players incurs linear regret on at least one of the instances. In particular, we take $\mathcal{I}$ and $\tilde{\mathcal{I}}$ to be the instances depicted in Table 1 with $\delta=O(1/\sqrt{T})$ (reproduced here for convenience).
@@
Fix any ALG for the centralized environment. Let $P$ (resp. $\tilde{P}$) denote the probability measure over canonical centralized bandit model when ALG is applied to the instance $\mathcal{I}$ (resp. $\tilde{\mathcal{I}}$) (see Section B.2). We will show that the regret is high on at least one of the instances $\mathcal{I}$ and $\tilde{\mathcal{I}}$.
@@
where $G^{c}$ is the complement of $G$. We apply Theorem B.2 to see that:
@@
where (1) applies Theorem B.2 and (2) applies Corollary B.4, and (3) uses the fact that $n_{T}(a_{1},b_{2})\leq T$. If we take $\delta=O(T^{-1/2})$, then we obtain a bound of $\Omega(T)$. Since these expected regrets are always nonnegative (see discussion above), this implies that either $R_{1}(T;\mathcal{I})=\Omega(T)$ or $R_{2}(T;\tilde{\mathcal{I}})=\Omega(T)$ as desired. ∎
B.5 Proof of Theorem 4.6
@@
Table 9: Hard instance for Theorem 4.6, where * is equal to $(0,0)$ for instance $\mathcal{I}_{a_{1},b_{1}}$, and $(0,2\delta)$ otherwise. Note that this example is structurally similar to the illustrative example in Table 3, but with $\left|\mathcal{A}\right|,\left|\mathcal{B}\right|\geq 2$.
@@
It suffices to prove this lower bound in a centralized environment where a single learner can choose action pairs $(a,b)$ and observes rewards for both players (Lemma B.1). We define a family of instances in the centralized game and evaluate the self-tolerant benchmarks on this family of instances. Arbitrarily pick some $(a_{1},b_{1})\in\mathcal{A}\times\mathcal{B}$ to be the “base” action. Let $\mathcal{F}_{\delta,\mathcal{A},\mathcal{B}}$ be the family of $|\mathcal{B}|$ instances of the form $(\mathcal{A},\mathcal{B},v_{1},v_{2})$ for varying settings of $v$, where we index the instances by $\mathcal{B}$. The utility functions for the instance $\mathcal{I}_{b^{\prime}}$ are equal to terms below (illustrated in Table 9):
@@
Fix an ALG be an algorithm for the centralized environment. For each $b\in\mathcal{B}$, let $P_{b}$ denote the probability measure over canonical centralized bandit model when ALG is applied to the instance $\mathcal{I}_{b}$ (see Section B.2). Let $n_{T}(a,b)=\sum_{t=1}^{T}\mathbbm{1}[a_{t}=a,b_{t}=b]$ be the random variable denoting the number of times that $(a,b)$ is pulled. We define:
@@
where $G^{c}$ is the complement of $G$. We apply Theorem B.2 to see that:
@@
where (1) applies Theorem B.2 and (2) applies Corollary B.4 and where (3) uses the fact that $\sum_{b\neq b_{1}}\sum_{a\neq a_{1}}\mathbb{E}_{P}[n_{T}(a,b)]\geq(|\mathcal{B}|-1)\sum_{a\neq a_{1}}\mathbb{E}_{P}[n_{T}(a,b_{m})]$.
@@
Appendix C Proofs for Section 4
C.1 Proof of Proposition 4.3